I had this question on my homework and I really don't understand it. Could you help explain how to solve this to me? Thank you!
"If K = {(x, y) | x - y = 5}, find the corresponding range of y for the domain x = 0, 2, 4"

Answer by ewatrrr(24785)

 
Hi
x - y = 5
0r y = x-5
x y
0 -5
2 -3
4 -1
y = -5, -3, -1